# The Ranking-Selberg integral on ${\bf GSp}(2)$ for square free levels

### v1

@@ 5.2. Computation of the average
for automorphic representations $\pi$ of ${\bf GL}_{2}({\mathbb{A}})$ ([24]) and the non-negativity of $L(\frac{1}{2},\pi_{i}\times\mathcal{AI}(\Lambda^{-1})=L(\frac{1}{2},{\rm BC}_{E/{\mathbb{Q}}}(\pi_{i})\otimes\Lambda^{-1})$ for $i=1,2$ ([16])).
Let us prove (5.16); by Theorem 5.3 (ii), we may assume $\lambda=(l,l)$. Suppose $\mu\not={\bf 1}$; then, by Theorem 5.3 (ii), we have that ${\mathbb{I}}^{(s)}((l,l),N,f_{S})^{{\rm SK,new}}\times s^{-1}$ is entire, hence $\mathbb{I}^{(0)}((l,l),N,f_{S})^{\rm SK,new}=0$ because $\widehat{L}(s,\mu)$, as well as $\widehat{L}(s,\pi_{0}\times\chi_{D})$, is entire. In the rest of the proof, we assume $\mu={\bf 1}$. Then, $s\widehat{L}(s,\mu)\widehat{L}(s+1,\mu)$ has a simple pole at $s=0$. By [38, Theorem 3.1 and Table 2], ${\rm SK}(\pi_{0})$ with $\pi_{0}\in\Pi_{{\rm cusp},{\bf PGL_{2}}}(2l-2,N)$ has the global $(T_{\theta},\Lambda)$-Bessel model only if the sign of the functional equation of $\pi_{0}$ is $-1$ so that $\widehat{L}(s+1/2,\pi_{0})$ has a zero at $s=0$. Hence, theorem 5.3 (ii) gives us the following majorant of ${\mathbb{I}}^{(0)}((l,l),N,f_{S})^{{\rm SK,new}}$.
$2^{3}\pi^{\frac{5}{2}}(l-1)^{-2}\frac{\Gamma(l)}{\Gamma(l-\frac{1}{2})}|D|^{\frac{l}{2}}[{\bf K}_{\rm{\bf{f}}}:{\bf K}_{0}(N)]^{-1}\sum_{\pi_{0}\in\Pi_{{\bf PGL}_{2},{\rm cusp}}^{(T_{\theta},{\bf 1})}(2l-2,N)}\frac{L(\frac{1}{2},\pi_{0}\times\chi_{D})\,|L^{\prime}(\frac{1}{2},\pi_{0})|}{L(1,\pi_{0};{\rm Ad})}.$
To estimate this, we invoke a subconvexity bound $L({1}/{2},\pi_{0}\times\chi_{D})=O((l^{2}N)^{\frac{1}{4}-\delta})$ for some $\delta>0$ from (5.17) and a lower bound $L(1,\pi_{0};{\rm Ad})>c_{0}\exp(-c_{1}\sqrt{\log(1+l^{2}N)})$ for some constant $c_{0},c_{1}>0$ which is known by [13, Theorem 0.1] (see also the remark after [22, Corollary 7]). The lower bound implies $L(1,\pi_{0};{\rm Ad})^{-1}=O((l^{2}N)^{2\delta})$. Moreover, by a common argument, a subconvexity bound for the central derivative of $L^{\prime}(1/2,\pi_{0})=O((l^{2}N)^{\frac{1}{4}-\delta})$ is derived from a subconvexity bound for $L(1/2,\pi_{0}|\cdot|_{\mathbb{A}}^{it})$ available from (5.17). By the uniform bound #$\#\Pi_{{\bf PGL}_{2},\rm cusp}(2l-2,N)=O(lN)$ and the asymptotic $\frac{\Gamma(l)}{\Gamma(l-\frac{1}{2})}\sim l^{\frac{1}{2}}\ (l\to\infty)$, the bound (5.16) follows. ∎
@@
Let $\pi_{i}$ ($i=1,2)$ be irreducible cuspidal automorphic representations of ${\bf PGL}_{n}$ such that $C(\pi_{1}),C(\pi_{2})\leqslant Q$ with $Q>2$. We assume that $\pi_{1}\not\cong\pi_{2}$ and both of them are self-dual and tempered everywhere. Then,
$L(1,\pi_{1}\times\pi_{2})\gg\exp(-C\sqrt{\log Q}).$
@@
is a Dirichlet series with non-negative coefficients ([14, Lemma a]). Moreover, by ([31, Proposition 9.4]) and due to $\pi_{1}$ and $\pi_{2}$ being self-dual, $L(s,\Pi\times\widetilde{\Pi})$ is expressed as
@@
which shows that $L(s,\Pi\times\widetilde{\Pi})$ has a pole of order $6$ at $s=1$ (due to $\pi_{1}\ncong\pi_{2}$), and has a zero of order $8$ at $s=\sigma$ if $L(\sigma+it,\pi_{i}\times\pi_{j})=0$ $(i,j=1,2)$. Hence, by [12, Lemma, p.178], one can show that $L(s,\pi_{1}\times\pi_{2})$ has no zeros in the interval $(1-\frac{C_{0}}{\log M},1)$ for some constant $C_{0}>0$ with $M=(1+|t|)^{24}C(\pi_{2}\times\pi_{2})^{18}C(\pi_{1}\times\pi_{1})^{9}C(\pi_{2}\times\pi_{2})^{2}$, where $C(\pi_{i}\times\pi_{j})$ is the analytic conductor of $L(s,\pi_{i}\times\pi_{j})$. By [14, lemma b], we have $C(\pi_{i}\times\pi_{j})\ll Q^{2n}$. Thus, for an absolute constant $C_{0}^{\prime}>0$, $L(s,\pi_{1}\times\pi_{2})$ has no zeros in the interval $(1-\frac{C_{0}^{\prime}}{\log Q},1)$. Now, we apply the argument of [22, Corollary 7]) to the $L$-function $L(s,\pi_{1}\times\pi_{2})$. Since $\pi_{1},\pi_{2}$ are assumed to be tempered everywhere, the optimal bound of the lambda function $\Lambda_{\mathcal{A}}(n)$ for $\mathcal{=}\pi_{1}\times\pi_{2}$ is available so that the automorphy of $\pi_{1}\times\pi_{2}$ is not necessary, which simplifies the proof to get the lower bound of $L(1,\pi_{1}\times\pi_{2})$. ∎
@@
(5.18) ${\mathbb{I}}^{(0)}(\lambda,N,f_{S})^{\rm{G,new}}={\mathbb{I}}^{(0)}(\lambda,N,f_{S})+O_{\Lambda,\mu,\lambda,f_{S}}(N^{-\frac{3}{2}}).$
@@ 6. Computation of local zeta-integrals for $p$-adic fields
We compute the local zeta-integral a la Piatetski-Shapiro ([30]) for several representations, taking particular test functions; as a result, we determine the local $L$-factors and the local $\varepsilon$-factors in [30] to confirm that they coincides with the expected ones listed in [35, Tables A8,A9]. As explained in §4.3, for a particular global application in mind, we only deal with representations of types I, IIb, IIIa and VIb (but allowing the central characters to be non-trivial.)
@@ 6.1. Local zeta integral for Bessel models
For a symmetric matrix $S=[\begin{smallmatrix}{\mathsf{a}}&\mathsf{b}/2\\ \mathsf{b}/2&\mathsf{c}\end{smallmatrix}]\in{\rm Sym}_{2}(F)$ such that ${\mathsf{d}}:=\mathsf{b}^{2}-4{\mathsf{a}}\mathsf{c}\neq 0$, set $\xi=[\begin{smallmatrix}\mathsf{b}/2&\mathsf{c}\\ -{\mathsf{a}}&-\mathsf{b}/2\end{smallmatrix}]$. Let $L:=F\oplus F\xi$ be the two-dimensional $F$-algebra, $\mathfrak{o}_{L}$ be its integer ring, and $\psi_{L}$ be the additive character on $L$ given by $\psi_{L}:=\psi\circ{\operatorname{tr}}_{L/F}$. The maps
@@
$L^{\times}=\left\{\tau\in{\bf GL}_{2}(F)\ \middle|\ {}^{t}\tau S\tau=\det\tau\cdot S\right\}.$
@@
$K_{0}^{\#}({\mathfrak{p}}^{e})$ $:=G^{\#}(F)\cap\left[\begin{smallmatrix}\mathfrak{o}_{L}&\mathfrak{o}_{L}\\ {\mathfrak{p}}^{e}\mathfrak{o}_{L}&\mathfrak{o}_{L}\end{smallmatrix}\right].$
@@
$\iota\left(\left[\begin{smallmatrix}\tau&0\\ 0&a\tau^{\dagger}\end{smallmatrix}\right]\right)=\left[\begin{smallmatrix}\tau&0\\ 0&a\cdot{}^{t}\tau^{\dagger}\end{smallmatrix}\right],\quad\tau\in L^{\times},a\in F^{\times},$
@@
the Bessel subgroup of $\mathsf{G}(F)$ (with respect to $S$). For a character $\Lambda:L^{\times}\to{\mathbb{C}}^{1}$, we can check that the map
$R\owns\iota\left(\left[\begin{smallmatrix}\tau&0\\ 0&\tau^{\dagger}\end{smallmatrix}\right]\right)\mathsf{n}(X)\mapsto\Lambda(\tau)\psi({\operatorname{tr}}{(SX)})\in{\mathbb{C}}^{1}$
defines a character on $R$. We denote this character by $\Lambda\otimes\psi_{S}$.
Let $(\pi,V_{\pi})$ be an irreducible admissible representation of $\mathsf{G}(F)$. Let $(V_{\pi}^{*})^{S,\Lambda}$ denote the space of all ${\mathbb{C}}$-linear forms $\ell:V_{\pi}\longrightarrow{\mathbb{C}}$ that satisfies
$\ell(\pi(r)\xi)=\Lambda\otimes\psi_{S}(r)\,\ell(\xi),\quad\xi\in V_{\pi},\,r\in R.$
Then, it holds that $\dim_{{\mathbb{C}}}(V_{\pi}^{*})^{S,\Lambda}\leqslant 1$ from ([27],[28]). We say that $\pi$ has a local $(S,\Lambda)$-Bessel model if $(V_{\pi}^{*})^{S,\Lambda}\not=(0)$. In this case, we can define the local $(S,\Lambda)$-Bessel model of $\pi$ as in §4.4. Moreover, when $\pi$ is spherical, we define the unramified $(S,\Lambda)$- Bessel datum $(\ell_{\pi}^{0},\xi_{\pi}^{0})\in(V_{\pi}^{*})^{S,\Lambda}\times V_{\pi}^{K}$ for $\pi$ as in §4.4, so that
@@
is the unramified Bessel function such that $B_{\pi}^{0}(1_{4})=1$. Recall that all irreducible admissible representations of $\mathsf{G}(F)$ that admit local Bessel models are classified in [35], and the result is conveniently summarized in [32, Table 2]. For our cases, as we mentioned in §4.4, we may assume that $\pi$ is of type I ,IIb, IIIa, or VIb. We refer some of its properties in ([32, Table 2]);
@@
Here, $\chi$, $\chi^{\prime}$, and $\sigma$ are unramified quasicharacters on $F^{\times}$ such that representations are irreducible and admit $(S,\Lambda)$-Bessel models. For $a\in F^{\times}$. For later use we set $\alpha=\chi(\varpi)$, $\beta=\chi^{\prime}(\varpi)$, and $\gamma=\sigma(\varpi)$. According to [35, Table A2], for type IIIa and type VIb, the corresponding representations are unitarizable if and only if the inducing quasi-characters are unitary, whereas, for type I, the unitarity of the inducing quasi-characters is equivalent to the corresponding representations being unitarizable and tempered. In what follows, having a global application in mind, we suppose the unitarity of all the inducing characters, i.e., $|\alpha|=|\beta|=|\gamma|=1$.
@@
where $\mu_{L}(\tau):=\mu(\tau\tau^{\dagger}),\tau\in L^{\times}$ and ${\rm{d}}^{\times}\tau$ is the normalized Haar measure on $L^{\times}$ such that ${\operatorname{vol}}(\mathfrak{o}_{L})=1$. For an irreducible smooth admissible representation $\pi$ of ${\mathsf{Z}}(F)\backslash\mathsf{G}(F)$, admitting $(S,\Lambda)$-Bessel model and $B\in{\mathscr{B}}(S,\Lambda)[\pi]$, define the zeta integral by
@@
It is shown that $Z(\phi,B,\mu;g)$ is a rational function in $q^{-s}$ ([41, Lemma 5.3.1]). Then, the local $L$-function $L(s,\pi,\mu)$ is defined to be the unique function such that $L(s,\pi,\mu)^{-1}$ is a polynomial in $q^{-s}$ with constant term $1$ and such that $Z(\phi,B,s,\mu,1)/L(s,\pi,\mu)\in{\mathbb{C}}[q^{s},q^{-s}]$ for all $\phi$ and $B$ ([30], see also [41, Proposition 5.3.3]). By [30, Proposition 3.2], there exists an entire function $\varepsilon(\pi,s,\mu,\psi)$ such that the local functional equation
@@
holds. Here, we set $B^{\dagger}\in{\mathscr{B}}(S,\Lambda^{\dagger})[\pi]$ as $B^{\dagger}(g):=B\left(\left[\begin{smallmatrix}1&{\mathsf{a}}^{-1}\mathsf{b}&0&0\\ 0&-1&0&0\\ 0&0&-1&0\\ 0&0&-{\mathsf{a}}^{-1}\mathsf{b}&1\end{smallmatrix}\right]g\right)$, $g\in\mathsf{G}(F)$, and $\Lambda^{\dagger}$ denote the Galois conjugate of $\Lambda$ defined by $\Lambda^{\dagger}(\tau):=\Lambda(\tau^{\dagger})$ for $\tau\in L^{\times}$, and $\widehat{\phi}$ is the Fourier transform of $\phi\in{\mathcal{S}}(L^{2})$ defined as
@@
with ${\rm{d}}u$ and ${\rm{d}}v$ being the Haar measures on $L$ such that ${\operatorname{vol}}(\mathfrak{o}_{L})=1$. The $L$-functions $L(s,\pi,\mu)$ are explicitly determined for several cases as shown in [41, Table 5], which, for $\mu=1$, coincide with the local $L$-functions listed in [35, Table A.8]:
(6.15) $\begin{array}[]{|c|c|c|}\hline\cr{{\rm type}}&{\pi}&{L(s,\pi)}\\ \hline\cr{{\rm I}}&{\chi\times\chi^{\prime}\rtimes\sigma}&{(1-\alpha\beta\gamma q^{-s})^{-1}(1-\alpha\gamma q^{-s})^{-1}(1-\beta\gamma q^{-s})^{-1}(1-\gamma q^{-s})^{-1}}\\ \hline\cr{{\rm IIb}}&{\chi 1_{{\bf GL}_{2}}\rtimes\sigma}&{(1-\alpha^{2}\gamma q^{-s})^{-1}(1-\gamma q^{-s})^{-1}(1-\alpha\gamma q^{-s-\frac{1}{2}})^{-1}(1-\alpha\gamma q^{-s+\frac{1}{2}})^{-1}}\\ \hline\cr{{\rm IIIa}}&{\chi\rtimes\sigma_{{\rm St}_{{\bf GL}_{2}}}}&{(1-\alpha\gamma q^{-s-\frac{1}{2}})^{-1}(1-\gamma q^{-s-\frac{1}{2}})^{-1}}\\ \hline\cr{{\rm VIb}}&{\tau(T,|\cdot|^{-\frac{1}{2}}\sigma)}&{(1-\gamma q^{-s-\frac{1}{2}})^{-2}}\\ \hline\cr\end{array}$

### v2

@@ 5.2. Computation of the average
for automorphic cuspidal representations $\pi$ of ${\bf GL}_{2}({\mathbb{A}})$ ([25]) and the non-negativity of $L(\frac{1}{2},\pi_{i}\times\mathcal{AI}(\Lambda^{-1})=L(\frac{1}{2},{\rm BC}_{E/{\mathbb{Q}}}(\pi_{i})\otimes\Lambda^{-1})$ for $i=1,2$ due to [16].
Let us prove (5.16); by Theorem 5.3 (ii), we may assume $\lambda=(l,l)$. Suppose $\mu\not={\bf 1}$; then, by Theorem 5.3 (ii), we have that ${\mathbb{I}}^{(s)}((l,l),N,f_{S})^{{\rm SK,new}}\times s^{-1}$ is entire, hence $\mathbb{I}^{(0)}((l,l),N,f_{S})^{\rm SK,new}=0$ because $\widehat{L}(s,\mu)$, as well as $\widehat{L}(s,\pi_{0}\times\chi_{D})$, is entire. In the rest of the proof, we assume $\mu={\bf 1}$. Then, $s\widehat{L}(s,\mu)\widehat{L}(s+1,\mu)$ has a simple pole at $s=0$. By [39, Theorem 3.1 and Table 2], ${\rm SK}(\pi_{0})$ with $\pi_{0}\in\Pi_{{\rm cusp},{\bf PGL_{2}}}(2l-2,N)$ has the global $(T_{\theta},\Lambda)$-Bessel model only if the sign of the functional equation of $\pi_{0}$ is $-1$ so that $\widehat{L}(s+1/2,\pi_{0})$ has a zero at $s=0$. Hence, theorem 5.3 (ii) gives us the following majorant of ${\mathbb{I}}^{(0)}((l,l),N,f_{S})^{{\rm SK,new}}$.
$2^{3}\pi^{\frac{5}{2}}(l-1)^{-2}\frac{\Gamma(l)\|f_{S}\|_{1}}{\Gamma(l-\frac{1}{2})}|D|^{\frac{l}{2}}[{\bf K}_{\rm{\bf{f}}}:{\bf K}_{0}(N)]^{-1}\sum_{\pi_{0}\in\Pi_{{\bf PGL}_{2},{\rm cusp}}^{(T_{\theta},{\bf 1})}(2l-2,N)}\frac{L(\frac{1}{2},\pi_{0}\times\chi_{D})\,|L^{\prime}(\frac{1}{2},\pi_{0})|}{L(1,\pi_{0};{\rm Ad})}.$
To estimate this, we invoke a subconvexity bound $L({1}/{2},\pi_{0}\times\chi_{D})=O((l^{2}N)^{\frac{1}{4}-\delta})$ for some $\delta>0$ from (5.17) and a lower bound $L(1,\pi_{0};{\rm Ad})>c_{0}\exp(-c_{1}\sqrt{\log(1+l^{2}N)})$ for some constant $c_{0},c_{1}>0$ which is known by [13, Theorem 0.1] (see also the remark after [23, Corollary 7]). The lower bound implies $L(1,\pi_{0};{\rm Ad})^{-1}=O((l^{2}N)^{2\delta})$. Moreover, by a common argument, a subconvexity bound for the central derivative $L^{\prime}(1/2,\pi_{0})=O((l^{2}N)^{\frac{1}{4}-\delta})$ is derived from the bound $L(1/2,\pi_{0}|\cdot|_{\mathbb{A}}^{it})=O((l^{2}N(1+|t|)^{1/4-\delta})$ that follows from (5.17). By the uniform bound #$\#\Pi_{{\bf PGL}_{2},\rm cusp}(2l-2,N)=O(lN)$ and the asymptotic $\frac{\Gamma(l)}{\Gamma(l-\frac{1}{2})}\sim l^{\frac{1}{2}}\ (l\to\infty)$, the bound (5.16) follows. ∎
@@
Let $\pi_{1}$ and $\pi_{2}$ be irreducible cuspidal automorphic representations of ${\bf PGL}_{n}$ such that $C(\pi_{1}),C(\pi_{2})\leqslant Q$ with $Q>2$. We assume that $\pi_{1}\not\cong\pi_{2}$ and both of them are self-dual and tempered everywhere. Then,
$L(1,\pi_{1}\times\pi_{2})\gg\exp(-C\sqrt{\log Q})$
@@
is a Dirichlet series with non-negative coefficients ([14, Lemma a]). Moreover, by ([32, Proposition 9.4]) and due to $\pi_{1}$ and $\pi_{2}$ being self-dual, $L(s,\Pi\times\widetilde{\Pi})$ is expressed as
@@
which shows that $L(s,\Pi\times\widetilde{\Pi})$ has a pole of order $6$ at $s=1$ (due to $\pi_{1}\ncong\pi_{2}$), and has a zero of order $8$ at $s=\sigma$ if $L(\sigma+it,\pi_{i}\times\pi_{j})=0$ $(i,j=1,2)$. Hence, by [12, Lemma, p.178], one can show that $L(s,\pi_{1}\times\pi_{2})$ has no zeros on the interval $(1-\frac{C_{0}}{\log M)},1)$ for some constant $C_{0}>0$ with $M=(1+|t|)^{24}C(\pi_{2}\times\pi_{2})^{18}C(\pi_{1}\times\pi_{1})^{9}C(\pi_{2}\times\pi_{2})^{2}$, where $C(\pi_{i}\times\pi_{j})$ is the analytic conductor of $L(s,\pi_{i}\times\pi_{j})$. By [14, lemma b], we have $C(\pi_{i}\times\pi_{j})\ll Q^{2n}$. Thus, for an absolute constant $C_{0}^{\prime}>0$, $L(s,\pi_{1}\times\pi_{2})$ is zero-free on the region $11-\frac{C_{0}^{\prime}}{\log Q(1+|{\operatorname{Im}}(s)|)}<{\operatorname{Re}}(s)<$ 1. Now, we apply the argument of [23, Corollary 7]) to the $L$-function $L(s,\pi_{1}\times\pi_{2})$. Since $\pi_{1},\pi_{2}$ are assumed to be tempered everywhere, the optimal bound of the lambda function $\Lambda_{\mathcal{A}}(n)$ for $\mathcal{A}=\pi_{1}\times\pi_{2}$ is available so that the automorphy of $\pi_{1}\times\pi_{2}$ is not necessary, which simplifies the proof to get the lower bound of $L(1,\pi_{1}\times\pi_{2})$. ∎
@@
(5.18) ${\mathbb{I}}^{(0)}(\lambda,N,f_{S})^{\rm{G,new}}={\mathbb{I}}^{(0)}(\lambda,N,f_{S})+O_{\Lambda,\lambda,\mu}(\|f_{S}\|_{1}N^{-\frac{3}{2}}).$
@@ 6. Computation of local zeta-integrals for $p$-adic fields
We compute the local zeta-integral a la Piatetski-Shapiro ([31]) for several representations, taking particular test functions; as a result, we determine the local $L$-factors and the local $\varepsilon$-factors in [31] to confirm that they coincides with the expected ones listed in [36, Tables A8,A9]. As explained in §4.3, for a particular global application in mind, we only deal with representations of types I, IIb, IIIa and VIb (but allowing the central characters to be non-trivial when we are concerned with newvectors.)
@@ 6.1. Local zeta integral for Bessel models
For a symmetric matrix $T=[\begin{smallmatrix}{\mathsf{a}}&\mathsf{b}/2\\ \mathsf{b}/2&\mathsf{c}\end{smallmatrix}]\in{\rm Sym}_{2}(F)$ such that ${\mathsf{d}}:=\mathsf{b}^{2}-4{\mathsf{a}}\mathsf{c}\neq 0$, set $\xi=[\begin{smallmatrix}\mathsf{b}/2&\mathsf{c}\\ -{\mathsf{a}}&-\mathsf{b}/2\end{smallmatrix}]$. Let $L:=F\oplus F\xi$ be the two-dimensional $F$-algebra, $\mathfrak{o}_{L}$ be its integer ring, and ${\mathfrak{p}}_{L}:={\mathfrak{p}}{\mathfrak{o}}_{L}$. Define the additive character $\psi_{L}$ on $L$ by $\psi_{L}:=\psi\circ{\operatorname{tr}}_{L/F}$. The maps
@@
$L^{\times}=\left\{\tau\in{\bf GL}_{2}(F)\ \middle|\ {}^{t}\tau T\tau=\det\tau\cdot T\right\}.$
@@
$K_{0}^{\#}({\mathfrak{p}}^{e})$ $:=G^{\#}(F)\cap\left[\begin{smallmatrix}\mathfrak{o}_{L}&\mathfrak{o}_{L}\\ {\mathfrak{p}}_{L}^{e}&\mathfrak{o}_{L}\end{smallmatrix}\right].$
@@
$\iota\left(\left[\begin{smallmatrix}\tau&0\\ 0&a\tau^{\dagger}\end{smallmatrix}\right]\right)=\mathsf{m}(\tau,a\det\tau),$
@@
the Bessel subgroup of $\mathsf{G}(F)$ (with respect to $T$). For a character $\Lambda:L^{\times}\to{\mathbb{C}}^{1}$, we can check that the map
$R\owns\iota\left(\left[\begin{smallmatrix}\tau&0\\ 0&\tau^{\dagger}\end{smallmatrix}\right]\right)\mathsf{n}(X)\mapsto\Lambda(\tau)\psi({\operatorname{tr}}{(TX)})\in{\mathbb{C}}^{1}$
defines a character on $R$. We denote this character by $\Lambda\otimes\psi_{T}$.
Let $(\pi,V_{\pi})$ be an irreducible admissible representation of $\mathsf{G}(F)$. Let $(V_{\pi}^{*})^{T,\Lambda}$ denote the space of all ${\mathbb{C}}$-linear forms $\ell:V_{\pi}\longrightarrow{\mathbb{C}}$ that satisfies
$\ell(\pi(r)\xi)=\Lambda\otimes\psi_{T}(r)\,\ell(\xi),\quad\xi\in V_{\pi},\,r\in R.$
Then, it holds that $\dim_{{\mathbb{C}}}(V_{\pi}^{*})^{T,\Lambda}\leqslant 1$ from ([28],[29]). We say that $\pi$ has a local $(T,\Lambda)$-Bessel model if $(V_{\pi}^{*})^{T,\Lambda}\not=(0)$. In this case, we can define the local $(T,\Lambda)$-Bessel model of $\pi$ as in §4.4. Moreover, when $\pi$ is spherical, we define the unramified $(T,\Lambda)$- Bessel datum $(\ell_{\pi}^{0},\xi_{\pi}^{0})\in(V_{\pi}^{*})^{T,\Lambda}\times V_{\pi}^{K}$ for $\pi$ as in §4.4, so that
@@
is the unramified Bessel function such that $B_{\pi}^{0}(1_{4})=1$. Recall that all irreducible admissible representations of $\mathsf{G}(F)$ that admit local Bessel models are classified in [36], and the result is conveniently summarized in [33, Table 2]. For our cases, as we mentioned in §4.4, we may assume that $\pi$ is of type I ,IIb, IIIa, or VIb. We refer some of its properties in ([33, Table 2]);
@@
Here, $\chi$, $\chi^{\prime}$, and $\sigma$ are unramified quasicharacters on $F^{\times}$ such that representations are irreducible and admit $(T,\Lambda)$-Bessel models. For $a\in F^{\times}$. For later use we set $\alpha=\chi(\varpi)$, $\beta=\chi^{\prime}(\varpi)$, and $\gamma=\sigma(\varpi)$. According to [36, Table A2], for type IIIa and type VIb, the corresponding representations are unitarizable if and only if the inducing quasi-characters are unitary, whereas, for type I, the unitarity of the inducing quasi-characters is equivalent to the corresponding representations being unitarizable and tempered. In what follows, having a global application in mind, we suppose the unitarity of all the inducing characters, i.e., $|\alpha|=|\beta|=|\gamma|=1$.
@@
where $\mu_{L}(\tau):=\mu(\tau\tau^{\dagger}),\tau\in L^{\times}$ and ${\rm{d}}^{\times}\tau$ is the normalized Haar measure on $L^{\times}$ such that ${\operatorname{vol}}(\mathfrak{o}_{L})=1$. For an irreducible smooth admissible representation $\pi$ of ${\mathsf{Z}}(F)\backslash\mathsf{G}(F)$, admitting $(T,\Lambda)$-Bessel model and $B\in{\mathscr{B}}(T,\Lambda)[\pi]$, define the zeta integral by
@@
It is shown that $Z(\phi,B,\mu;g)$ is a rational function in $q^{-s}$ ([42, Lemma 5.3.1]). Then, the local $L$-function $L(s,\pi,\mu)$ is defined to be the unique function such that $L(s,\pi,\mu)^{-1}$ is a polynomial in $q^{-s}$ with constant term $1$ and such that $\tfrac{Z(\phi,B,s,\mu,1)}{L(s,\pi,\mu)}\in{\mathbb{C}}[q^{s},q^{-s}]$ for all $\phi$ and $B$ ([31], see also [42, Proposition 5.3.3]). By [31, Proposition 3.2], there exists an entire function $\varepsilon(\pi,s,\mu,\psi)$ such that the local functional equation
@@
holds. Here, we set $B^{\dagger}\in{\mathscr{B}}(T,\Lambda^{\dagger})[\pi]$ as $B^{\dagger}(g):=B\left(\left[\begin{smallmatrix}1&{\mathsf{a}}^{-1}\mathsf{b}&0&0\\ 0&-1&0&0\\ 0&0&-1&0\\ 0&0&-{\mathsf{a}}^{-1}\mathsf{b}&1\end{smallmatrix}\right]g\right)$, $g\in\mathsf{G}(F)$, and $\Lambda^{\dagger}$ denote the Galois conjugate of $\Lambda$ defined by $\Lambda^{\dagger}(\tau):=\Lambda(\tau^{\dagger})$ for $\tau\in L^{\times}$, and $\widehat{\phi}$ is the Fourier transform of $\phi\in{\mathcal{S}}(L^{2})$ defined as
@@
with ${\rm{d}}u$ and ${\rm{d}}v$ being the Haar measures on $L$ such that ${\operatorname{vol}}(\mathfrak{o}_{L})=1$.